OpenStudy (cameron2000):

Simplify (4x − 6) + (3x + 6). 7x 7x − 12 x 7x + 12

OpenStudy (skullpatrol):

any ideas?

OpenStudy (cameron2000):

I think it is A but i dont know

OpenStudy (skullpatrol):

(4x - 6) + (3x + 6) = 4x - 6 + 3x + 6 = 4x + 3x - 6 + 6 = 7x

OpenStudy (skullpatrol):

A is correct :D

OpenStudy (cameron2000):

ok thank you

OpenStudy (skullpatrol):

Thanks for trying to learn!